# Turing’s Test, a Beautiful Thought Experiment

### v2

@@ 4. The method of thought experiments
To understand this better, let us take a brief look at Turing’s 1948 report “Intelligent Machinery” (op. cit.). In section (§3) ‘Varieties of machinery,’ he noted: “All machinery can be regarded as continuous, but when it is possible to regard it as discrete it is usually best to do so.” A brain, he wrote, “is probably” a ‘continuous controlling’ machine, but in light of the digital nature of neural impulses, it “is very similar to much discrete machinery.” In section (§6) ‘Man as Machine,’ he referred to the imitation of “any small part of a man” by machines: “A great positive reason for believing in the possibility of making thinking machinery is the fact that it is possible to make machinery to imitate any small part of a man” (p. 420). In light of this, he argued: “One way of setting about our task of building a ‘thinking machine’ would be to take a man as a whole and to try to replace all the parts of him by machinery.” But Turing dismissed such a method as “altogether too slow and impracticable,” and later alluded to moral and aesthetic reasons as well. [^16] For the May 1951 broadcast (op. cit.), he wrote: “I certainly hope and believe that no great efforts will be put into making machines with the most distinctively human, but non-intellectual characteristics such as the shape of the human body; it appears to me to be quite futile to make such attempts and their results would have something like the unpleasant quality of artificial flowers.”
@@
We can now follow Turing’s use of the method of continuous variation in the design of his imitation tests. The essential question ($Q^{\star}$) Turing asks is whether the intellectual and cultural performances (the stereotypes) [^17] S. Sterrett first emphasized the importance of stereotypes in the imitation game (Sterrett, 2000). associated with woman, man, machine (the types) could be imitated, and thus softly transposed. Note that for any arbitrarily chosen type, say, a woman, further specific subtypes can be continuously conceived and considered as varied conditions of the imitation game: women having property $p$, women having subproperty $p^{\prime}\subset p$, and so on. For any two arbitrarily chosen types, a new type can be conceived, whether as a specialization or a modification (“any small part of a man”). Because concepts are fluid entities, there is an evolving continuum of levels and types.
@@
The question across the various versions of the game can be posed this way: how does C’s perception of A’s performance against B’s performance change as the game’s conditions are (continuously) varied? Will it change if gendered verbal behavior is required as a subtype of human verbal behavior? Will it change if the machine’s hardware is increased and/or its learning program is modified? For Turing, there is no conceptual discontinuity among the various conditions that instantiate his thought experiment.
@@
From 1948 to 1952, Turing presented various imitation tests based on both the game of chess and conversation. A historically sound problem, because it does not struggle with the materiality of Turing’s texts and their chronological coherence, does not erase some of his tests in favor of others, does not ignore the historical conditions of his proposal, is ($P_{2}$) Why would Turing present multiple versions of his test, as opposed to a well-defined, controlled experiment? I have presented an answer by reconstructing Turing’s use of “the method of thought experiments,” and this will be reinforced in the next section. Turing’s purpose in his 1950 paper was not to propose a ‘benchmark’ for the non-existent field of AI, but to respond to critics. Especially in 1949, he felt compelled to point out that the new science of universal digital computing would eventually have an impact and expand our view of ‘thinking.’
5. 1949, the crucial year
@@
As is often the case with thought experiments, Turing proposed his test out of a controversy (Gonçalves, 2023d). He was coming from his continuing disputes with the physicist and computer pioneer, Fellow of the Royal Society (FRS), Douglas Hartree (1897-1958), over the meaning of the newly existing digital computers, which had started in 1946 (Gonçalves, 2023c). Now, in mid-1949, new opponents had arrived, notably the neurosurgeon Geoffrey Jefferson (1886-1961), and the chemist and philosopher Michael Polanyi (1891-1976), both also FRS and based at the same institution as Turing, the University of Manchester, where Turing had spent a year as a Reader in the Department of Mathematics (Hodges, 1983). These three thinkers challenged Turing’s claims about the future possibilities and limitations of digital computers.
@@
In June 1949, Hartree published his Calculating Instruments and Machines (Hartree, 1949), in which Ada Lovelace’s work was acknowledged seemingly for the first time by a twentieth-century computer pioneer (Gonçalves, 2023c). Since November 1946, Hartree had been opposing the use of the term ‘electronic brain.’ He wrote in a letter to the Times: “These machines can only do precisely what they are instructed to do by the operators who set them up.”181818“The ‘Electronic Brain’: A Misleading Term; No Substitute for Thought,” Times, November 7, 1946. Now in 1949, Hartree added strength to his argument by quoting the words of Ada Lovelace from the 1840s about Charles Babbage’s machine: “The Analytical Engine has no pretensions to originate anything $\ldots$ It can do whatever we know how to order it to perform” (her emphasis) (Hartree, 1949, p. 70). Noting Hartree’s anachronism in taking Lovelace’s words out of their time and place, Turing further developed his earlier, 1947 response to Hartree’s challenge,191919‘Lecture to L.M.S. Feb. 20 1947’, p. 22; op. cit. now calling it ‘(6) Lady Lovelace’s objection’ (Turing, 1950, p. 450). Turing argued that intelligent behavior is the result of learning, a capability he had no problem attributing to future digital computers. He also questioned the implicit assumption of Hartree’s challenge: “Who can be certain that ‘original work’ that he has done was not simply the growth of the seed planted in him by teaching, or the effect of following well-known general principles” (p. 450). In the imitation game, Turing suggested, the interrogator would be able to evaluate the machine’s ability to learn: “The game (with the player B omitted) is frequently used in practice under the name of viva voce to discover whether some one really understands something or has ‘learnt it parrot fashion’ ” (p. 446). But then we might ask, what is player B doing in the imitation game? Following the 1949 events will suggest an answer.
@@
On June 9, in London, Jefferson delivered his prestigious Lister Oration on ‘The Mind of Mechanical Man,’ which was published in the debuting British Medical Journal on June 25 (Jefferson, 1949). His lecture was headlined in the Times on June 10,202020‘No Mind For Mechanical Man.’ Times, 10 June 1949, p. 2. emphasizing his claim that “Not until a machine can write a sonnet or compose a concerto because of thoughts and emotions felt, and not by the chance fall of symbols, could we agree that machine equals brain” (p. 1110). This rendered Turing’s ironic response: “I do not think you can even draw the line about sonnets, though the comparison is perhaps a little bit unfair because a sonnet written by a machine will be better appreciated by another machine.”212121‘The Mechanical Brain.’ Times, 11 June 1949, p. 4. In October and December 1949, two seminars on ‘Mind and Machine’ were organized by Polanyi et al., and attended by Jefferson, Turing et al., at the Philosophy Department in Manchester (Polanyi, 1958, p. 275; cf. also note 1 above). These seminar discussions, followed by Jefferson giving Turing an offprint of his Lister Oration, [^22] This may have happened in the evening of the December meeting of the Manchester seminar (op. cit.), when, according to a later letter from Jefferson to Ethel S. Turing, Turing and J.Z. Young went to dinner at Jefferson’s house (Turing, 1959, p. xx). which Turing read and marked with a pencil, [^23] Off-print, ‘The mind of mechanical man’ by Geoffrey Jefferson. Archives Centre, King’s College, Cambridge, AMT/B/44. led him to write his 1950 paper and propose his test (Gonçalves, 2023d).
@@
In his Lister Oration (Jefferson, 1949), Jefferson had characterized intelligence as an emergent property of the animal nervous system. He emphasized that “sex hormones introduce peculiarities of behaviour often as inexplicable as they are impressive” (p. 1107). Because “modern automata” are not moved by male and female sex hormones, they could not exhibit such peculiarities to imitate the actions of animals or ‘men.’ Specifically, he used a thought experiment to criticize Grey Walter’s mechanical turtles by suggesting that gendered behavior is causally related to the physiology of sex hormones (ibid.):
@@
Jefferson thus brought forward the image of a genuine individual of a kind, which is placed side by side with the artificial one so that the latter’s artificiality is emphasized. The function of the genuine individual is to expose the artificiality of the impostor. The means of exposure is to fail at demonstrating interesting (sexual) behavior. This can explain Turing’s introduction of a (gendered) control player B, who appears in Turing’s 1950 test, whose design was prompted by his reading of Jefferson, but not in Turing’s 1948, 1951, and 1952 tests. In discussing “(4) The Argument from Consciousness,” Turing addressed Jefferson directly and quoted in full his conditions for agreeing “that machine equals brain,” including “be warmed by flattery” and “be charmed by sex” (Turing, 1950, pp. 445-446). In discussing the “(5) Argument from Various Disabilities,” Turing again mentioned Jefferson and argued that to say that a machine could never “fall in love” or “make someone fall in love with it” was a flawed scientific induction from the capabilities of present machines.
Thus, an answer to the first part of problem $P_{3}$, why gender imitation in a test of machine intelligence, is that Turing’s design of his test was an ironic response to Jefferson’s association of sex and gender with intelligence, particularly his suggestion that gendered behavior is causally related to the physiology of male and female sex hormones. It remains to address the second part of problem $P_{3}$, why Turing chose conversation as the intelligence task of his test.
@@
Surviving minutes of the ‘Mind and Machine’ seminar held on October 27, 1949, were published in 2000 by a participant, Wolfe Mays. [^24] W. Mays, ‘Turing and Polanyi on minds and machines.’ Appraisal, 3(2), 55-62. Andrew Hodges also published it on his website: https://www.turing.org.uk/sources/wmays1.html. Accessed December 6, 2023. In the first session, Polanyi presented a statement, ‘Can the mind be represented by a machine?,’ [^25] Polanyi, Michael. Papers, Box 22, Folder 19, Hanna Holborn Gray Special Collections Research Center, University of Chicago Library. which was a Gödelian argument that humans can do things that machines cannot. Although Turing had already addressed this argument in his 1947 lecture (op. cit.), Polanyi’s insistence may help explain Turing’s inclusion of “(3) The Mathematical Objection” (Turing, 1950). Further, the minutes (op. cit.) show that Polanyi tried to distinguish the formal “rules of the logical system” from the informal “rules which determine our own behaviour,” and this can explain Turing’s inclusion of “(8) The Argument from Informality of Behaviour” (Turing, 1950).
@@
Years later (Polanyi, 1958, p. 275), Polanyi remembered “a communication to a Symposium held on ‘Mind and Machine’ at Manchester University in October, 1949,” in which “A.M. Turing has shown that it is possible to devise a machine which will both construct and assert as new axioms an indefinite sequence of Gödelian sentences.” [^26] Polanyi added that “this is foreshadowed” in Turing’s 1938 paper based on his Ph.D. thesis, ‘Systems of Logic Based on Ordinals,’ J. London Math. Soc. s2-45(1), 161-228. Polanyi resumed, showing that he assimilated the punch: “Any heuristic process of a routine character—for which in the deductive sciences the Gödelian process is an example—could likewise be carried out automatically.” However, Polanyi used the same argument to dismiss the game of chess as a testbed for machine intelligence, noting: “A routine game of chess can be played automatically by a machine, and indeed, all arts can be performed automatically to the extent to which the rules of the art can be specified.”
@@
Chess, not conversation, had been Turing’s chosen field to illustrate, develop, and test machine intelligence since at least February 1946.272727‘Proposed electronic calculator,’ February 1946. Archives Centre, King’s College, Cambridge, AMT/C/32. On p. 16, Turing asks: “Can the machine play chess?” In his 1948 ‘Intelligent Machinery’ report (op. cit.), Turing had discussed a tradeoff between convenient and impressive intellectual fields for exploring machine intelligence. After discussing “various games e.g. chess,” Turing wrote: “Of the above possible fields the learning of languages would be the most impressive, since it is the most human of these activities.” However, he avoided language learning because it seemed “to depend rather too much on sense organs and locomotion to be feasible,” stuck with chess, and ended up describing a chess-based imitation game. Now in October 1949, he saw chess being dismissed as an unimpressive to make the case for machine intelligence because its rules could be specified.
@@
Some time later, probably around Christmas 1949, Turing will read Jefferson’s Lister Oration (Jefferson, 1949) and mark the passage quoting René Descartes (p. 1106), which starts: “Descartes made the point, and a basic one it is, that a parrot repeated only what it had been taught and only a fragment of that; it never used words to express its own thoughts.” Overall, Jefferson suggested ‘speech’ to be the distinguishing feature of human intelligence compared to other kinds of animal intelligence: “Granted that much that goes on in our heads is wordless $\ldots$ we certainly require words for conceptual thinking as well as for expression $\ldots$ It is here that there is the sudden and mysterious leap from the highest animal to man, and it is in the speech areas of the dominant hemisphere $\ldots$ that Descartes should have put the soul, the highest intellectual faculties” (p. 1109).
@@
Unlike chess, which is governed by definite rules, good performance in conversation was thought to be reserved for humans. So Turing’s 1950 change to “the learning of languages” as the intellectual field addressed by his test can best be understood as yet another concession to Jefferson, and in this case to Polanyi as well.
In summary, Turing varied the design of his imitation tests to respond to the challenges posed by Hartree, Polanyi, and Jefferson. This fits neatly into Popper’s methodological rule for “the use of imaginary experiments in critical argumentation” (Popper, 1959): “the idealizations made must be concessions to the opponent, or at least acceptable to the opponent” (p. 466, Popper’s emphasis).
6. A hammer and a feather on the moon
On August 2, 1971, more than three centuries after Galileo’s death, a live anecdotal demonstration of Galileo’s legendary tower experiment was performed for the television cameras by astronaut David Scott during the final Apollo 15 moonwalk (Fig. 3). Far from the Earth’s atmosphere, essentially in a vacuum, the astronaut simultaneously released a heavy object (an aluminum geological hammer) and a light object (a falcon feather) from approximately the same height, which fell to the ground at the same rate to the naked eye. The performer, who attributed their successful mission in part to “a rather significant discovery about falling objects in gravity fields” made long ago by “a gentleman named Galileo,” celebrated: “How about that! Mr Galileo was correct in his findings.” [^28] For footage and a technical description of the demonstration, see ¡http://nssdc.gsfc.nasa.gov/planetary/lunar/apollo_15_feather_drop.html¿. Accessed December 6, 2023.
Variants of Galileo’s falling-bodies experiment first appeared in his De motu drafts written in the 1590s. Decades later came the punchy presentation of his 1638 Two New Sciences (Galilei, 1638, pp. 66-67):
SALVIATI: But without experiences, by a short and conclusive demonstration, we can prove clearly that it is not true that a heavier moveable is moved more swiftly than another, less heavy, these being of the same material, and in a word, those of which Aristotle speaks. Tell me, Simplicio, whether you assume that for every heavy falling body there is a speed determined by nature such that this cannot be increased or diminished except by using force or opposing some impediment to it $\ldots$ [SIMPLICIO agrees]
Then if we had two moveables whose natural speeds were unequal, it is evident that were we to connect the slower to the faster, the latter would be partly retarded by the slower, and this would be partly speeded up by the faster $\ldots$
Figure 3. Alan Bean’s painting “The Hammer and the Feather,” 1986. Reproduced with permission. Source: http://alanbeangallery.com/hammerfeather-story.html.
From the outset, Galileo makes Simplicio accept Salviati’s carefully formulated assumption that for every heavy falling body there is a natural speed that cannot be altered except by external intervention. However, the Aristotelian could have found a way out by noting the imprecision and denying that weight and natural speed are physically determinate for connected but not unified bodies (Gendler, 1998).
Could Galileo have run a variant of this experiment from the Leaning Tower of Pisa and obtained the results claimed in the story? Physicists went to the laboratory and concluded that it is doubtful (Adler and Coulter, 1978). Decades earlier, it was found (Cooper, 1935) that the only source for the legendary story was actually an apologetic biography written by Galileo’s disciple Vincenzo Viviani 12 years after Galileo’s death and first published in 1717. For centuries, Galileo was largely considered “the first true empiricist,” but the role of experiment in Galilean science was more complex than previously thought (Segre, 1989; Palmieri, 2005).
Galileo’s falling-bodies experiment suggests an anomaly in Aristotle’s theory of motion under certain idealized conditions, the existence of motion in a void, which was unacceptable at the time. Testing such an existential hypothesis was in infinite regress with the conditions it required, and creating those conditions would require long-term scientific and technological progress.
The Galilean impasse could only be broken by the power of his thought experiments as propaganda292929‘Propaganda’ is meant here in its pre-Nazi, neutral sense of propagating, spreading. for a next generation of scientists (Gonçalves, 2023a). Robert Boyle (1627-1692) was one of them. As a teenager, he visited Florence shortly before Galileo’s death and was impressed by “the new paradoxes of the great star-gazer Galileo” (Fulton, 1960, p. 119). In his career, Boyle built air pumps and special chambers to study vertical fall in small evacuated environments, and became an exponent of experimental philosophy in the Royal Society. There is no record of Boyle interpreting Galilean science literally and performing any tower experiments. The value of Galileo’s thought experiments was to lay conceptual foundations and to conjecture a class of idealized phenomena to be pursued by progressive science and technology. There is a path that connects Galileo’s thought experiments and Boyle’s vacuum chambers to the space programs of the 1950s and finally to the anecdotal confirmation of Galileo’s hypothesis by the crew of the Apollo 15 mission to the moon.
Now note the analogy with the Turing test (Gonçalves, 2023a). As Hayes and Ford claimed: “The tests are circular: they define the qualities they are claiming to be evidence for” (p. 974). Turing’s existential hypothesis of a Turing-test-passing machine is in infinite regress with the conditions assumed by the test: an idealized computer equipped with a hypothetically appropriate program (Box 1). But if these conditions exist, why do we need a Turing test at all? This shows that practical Turing tests can serve at best as anecdotal confirmation of Turing’s hypothesis, and at worst, as we have seen for decades, as publicity stunts. The value of Turing’s test must lie elsewhere.
7. The value of Turing’s test for AI

### v3

@@ Abstract
To understand this better, let us take a brief look at Turing’s 1948 report “Intelligent Machinery” (op. cit.). In section (§3) ‘Varieties of machinery,’ he noted: “All machinery can be regarded as continuous, but when it is possible to regard it as discrete it is usually best to do so.” A brain, he wrote, “is probably” a ‘continuous controlling’ machine, but in light of the digital nature of neural impulses, it “is very similar to much discrete machinery.” In section (§6) “Man as Machine,” he referred to the imitation of “any small part of a man” by machines: “A great positive reason for believing in the possibility of making thinking machinery is the fact that it is possible to make machinery to imitate any small part of a man” (p. 420). In light of this, he argued: “One way of setting about our task of building a ‘thinking machine’ would be to take a man as a whole and to try to replace all the parts of him by machinery.” But Turing dismissed such a method as “altogether too slow and impracticable,” and later alluded to moral and aesthetic reasons as well. [^18] For the May 1951 broadcast (op. cit.), he wrote: “I certainly hope and believe that no great efforts will be put into making machines with the most distinctively human, but non-intellectual characteristics such as the shape of the human body; it appears to me to be quite futile to make such attempts and their results would have something like the unpleasant quality of artificial flowers.”
@@
We can now follow Turing’s use of the method of continuous variation in the design of his imitation tests. The central question ($Q^{\star}$) Turing asks is whether the intellectual and cultural performances (the stereotypes) [^19] Susan Sterrett first emphasized the importance of stereotypes in the imitation game [sterrett2000]. associated with woman, man, machine (the types) could be imitated, and thus softly transposed. Note that for any arbitrarily chosen type, say, a woman, further specific subtypes can be continuously conceived and considered as varied conditions of the imitation game: women having a certain property, a subproperty, and so on. For any two arbitrarily chosen types, a new type can be conceived, whether as a specialization or a modification (“any small part of a man”). Because concepts are fluid abstractions, there is an evolving continuum of levels and types.
@@
The question across the various versions of the game can be posed this way: how does C’s perception of A’s performance against B’s performance change as the game’s conditions are (continuously) varied? Will it change if gendered verbal behavior is required as a subtype of human verbal behavior? Will it change if the machine’s hardware is increased and/or its learning program is modified? For Turing, there is no conceptual discontinuity between the various conditions instantiating his thought experiment. It is often asked whether the imitation goal changes from the machine-woman test (p. 434) to the machine-man test (p. 442). Note that this open question vanishes under this interpretation of the test, which observes material and structural facts of Turing’s text: he describes such goals only very loosely, and, more importantly, continuously varies the design of the game. His focus is on the power of universal digital computing to imitate and deconstruct arbitrary types, not on setting this or that type for the sake of promoting a particular test.
@@
From 1948 to 1952, Turing presented various imitation tests based on both the game of chess and conversation, even bringing back chess at the end of his 1950 paper (p. 460), and referring to his (“my”) various “imitation tests” in 1952.202020‘Can automatic calculating machines be said to think?’, January 1952 (op. cit.). Why would he present various tests, as opposed to a well-defined, controlled experiment? This is a historically sound question, because it does not struggle with the materiality of Turing’s texts and their chronological coherence. Nor does it erase some of his tests in favor of others, or overlook the historical conditions of his proposal. This paper has provided an answer by reconstructing what can be called Turing’s use of the method of thought experiments, whose context will now be explored.
1949, THE CRUCIAL YEAR
@@
As is often the case with thought experiments, Turing proposed his test out of a controversy [goncalves2023argument]. He was coming from his continuing disputes with the physicist and computer pioneer, Fellow of the Royal Society (FRS), Douglas Hartree (1897-1958), over the meaning of the newly existing digital computers, which had started in 1946 [goncalves2023lovelace]. Now, in mid-1949, new opponents had arrived, most notably the neurosurgeon Geoffrey Jefferson (1886-1961), and the chemist and philosopher Michael Polanyi (1891-1976), both also FRS and based at the same institution as Turing, the University of Manchester, where Turing had spent a year as a Reader in the Department of Mathematics [hodges1983]. These three thinkers challenged Turing’s claims about the future possibilities and limitations of digital computers.
@@
In June 1949, Hartree published his Calculating Instruments and Machines [hartree1949], in which Ada Lovelace’s work was acknowledged seemingly for the first time by a twentieth-century computer pioneer [goncalves2023lovelace]. Since November 1946, Hartree had been opposing the use of the term “electronic brain.” He wrote in a letter to the Times: “These machines can only do precisely what they are instructed to do by the operators who set them up.”212121“The ‘Electronic Brain’: A Misleading Term; No Substitute for Thought,” Times, November 7, 1946. Now in 1949, Hartree added strength to his argument by quoting the words of Ada Lovelace from the 1840s about Charles Babbage’s machine: “The Analytical Engine has no pretensions to originate anything $\ldots$ It can do whatever we know how to order it to perform” (her emphasis) [hartree1949, p. 70]. Noting Hartree’s anachronism in taking Lovelace’s words out of their time and place, Turing further developed his earlier, 1947 response to Hartree’s challenge,222222‘Lecture to L.M.S. Feb. 20 1947’ (op. cit.), p. 22. now calling it “(6) Lady Lovelace’s objection” [turing1950, p. 450]. Turing argued that intelligent behavior is the result of learning, a capability he had no problem attributing to future digital computers. He also questioned the implicit assumption of Hartree’s challenge: “Who can be certain that ‘original work’ that he has done was not simply the growth of the seed planted in him by teaching, or the effect of following well-known general principles” (p. 450). In the imitation game, Turing suggested, the interrogator would be able to evaluate the machine’s ability to learn: “The game (with the player B omitted) is frequently used in practice under the name of viva voce to discover whether some one really understands something or has ‘learnt it parrot fashion’ ” (p. 446). But then, what is player B doing in the imitation game? Following the 1949 events will suggest an answer.
@@
On June 9, in London, Jefferson delivered his prestigious Lister Oration on “The Mind of Mechanical Man,” which was published in the debuting British Medical Journal on June 25 [jefferson1949]. His lecture was headlined in the Times on June 10,232323“No Mind For Mechanical Man.” Times, 10 June 1949, p. 2. emphasizing his claim that “Not until a machine can write a sonnet or compose a concerto because of thoughts and emotions felt, and not by the chance fall of symbols, could we agree that machine equals brain” (p. 1110). This rendered Turing’s famous response: “I do not think you can even draw the line about sonnets, though the comparison is perhaps a little bit unfair because a sonnet written by a machine will be better appreciated by another machine.”242424“The Mechanical Brain.” Times, 11 June 1949, p. 4. In October and December 1949, the two seminars on “Mind and Machine” were organized by Polanyi et al., and attended by Jefferson, Turing et al., at the Philosophy Department in Manchester [[]p. 275]polanyi1958. These seminar discussions, followed by Jefferson giving Turing an offprint of his Lister Oration, [^25] This may have happened in the evening of the December meeting of the Manchester seminar (op. cit.), when, according to a later letter from Jefferson to Ethel S. Turing, Turing and J.Z. Young went to dinner at Jefferson’s house [sara1959, p. xx]. which Turing read and marked with a pencil, [^26] Off-print, “The mind of mechanical man” by Geoffrey Jefferson. Archives Centre, King’s College, Cambridge, AMT/B/44. led him to write his 1950 paper and propose his test [goncalves2023argument].
@@
Jefferson had characterized intelligence as an emergent property of the animal nervous system. He emphasized that “sex hormones introduce peculiarities of behaviour often as inexplicable as they are impressive” (p. 1107). Because “modern automata” are not moved by male and female sex hormones, they could not exhibit such peculiarities to imitate the actions of animals or “men.” Specifically, he used a thought experiment to criticize Grey Walter’s mechanical turtles by suggesting that gendered behavior is causally related to the physiology of sex hormones (ibid.):
@@
In reaction to Grey Walter and his transgressive tortoises [hayward2001], [^27] Jefferson would attack Walter’s automata again in speeches to learned societies in Manchester in 1953 [jefferson1953] and 1956 [jefferson1956] in the wake of Walter’s The Living Brain [walter1953]. Jefferson offered the image of a genuine individual of a species placed side by side with an artificial one to emphasize the latter’s artificiality. The function of the genuine individual is to expose the artificiality of the impostor. The means of exposure is the failure to demonstrate the authentic (sexual) behavior of the original species. This can explain Turing’s introduction of a (gendered) control player B, who appears in Turing’s 1950 test, whose design was prompted by his reading of Jefferson, but not in Turing’s 1948, 1951, and 1952 tests. In discussing “(4) The Argument from Consciousness,” Turing addressed Jefferson directly and quoted in full his conditions for agreeing “that machine equals brain,” including “be warmed by flattery” and “be charmed by sex” [turing1950, pp. 445-446]. [^28] Jefferson’s response was [jefferson1953, p. 73]: “But there are those like Dr. Turing who believe that we have no right to deny self-consciousness to the machines since they fulfill the definition of mind as given above — the ability to make choices.” In discussing the “(5) Argument from Various Disabilities” (p. 447), Turing again addressed Jefferson (p. 450) and argued that to say that a machine could never “fall in love” or “make someone fall in love with it” was a flawed scientific induction from the capabilities of present machines.
Turing’s test design may have been an ironic response to Jefferson’s suggestion that gendered behavior is causally related to the physiology of male and female sex hormones. As a repressed homosexual [hodges1983] and non-conformist in postwar England [goncalves2023irony], Turing might have been prepared to refer to a gender test. However, we have just seen that a basic version of this idea was actually available to him in the form of Jefferson’s reaction to Walter’s tortoises. Apart from Turing’s 1950 paper, in which he is in direct dialogue with Jefferson, he links his views on machine intelligence to sex and gender in one other known source, again with Jefferson in the background. At the end of a letter to a friend written after the Wilmslow police challenged his testimony, [^29] Turing to Norman Routledge, circa mid-Feb., 1952. Archives Centre, King’s College, Cambridge, AMT/D/14a. he comments on the BBC broadcast of January 1952 (op. cit.): “Glad you enjoyed the broadcast. J. [Jefferson] certainly was rather disappointing though. I’m rather afraid that the following syllogism may be used by some in the future[:] Turing believes machines think. Turing lies with men. Therefore machines do not think.” It remains to be explored Turing’s choice of conversation for his test.
@@
Surviving minutes of the “Mind and Machine” seminar held on October 27, 1949, were published in 2000 by a participant, Wolfe Mays [mays2000]. In the first session, Polanyi presented a statement, “Can the mind be represented by a machine?,” [^30] Polanyi, Michael. Papers, Box 32, Folder 6, Hanna Holborn Gray Special Collections Research Center, University of Chicago Library. which was a Gödelian argument that humans can do things that machines cannot. Although Turing had already addressed this argument in his 1947 lecture (op. cit.), Polanyi’s insistence may help explain Turing’s inclusion of “(3) The Mathematical Objection” [turing1950, p. 444]. Further, the minutes of the Manchester seminar show that Polanyi tried to distinguish the formal “rules of the logical system” from the informal “rules which determine our own behaviour,” and this helps explain Turing’s inclusion of “(8) The Argument from Informality of Behaviour” (p. 452). Polanyi’s argument came too late, as Turing had long been involved in such debates in the Moral Sciences Club at Cambridge University, both before and after World War II. [^31] Minutes and other papers of the Moral Sciences Club, 1878–2018, Cambridge University Library, GBR/0265/UA/Min.IX.39-48, 56-6 etc.
@@
Years later [polanyi1958, p. 275], Polanyi remembered “a communication to a Symposium held on ‘Mind and Machine’ at Manchester University in October, 1949,” in which “A.M. Turing has shown that it is possible to devise a machine which will both construct and assert as new axioms an indefinite sequence of Gödelian sentences.” [^32] Polanyi added that “this is foreshadowed” in Turing’s 1938 paper based on his Ph.D. thesis, “Systems of Logic Based on Ordinals,” J. London Math. Soc. s2-45(1), 161-228. Polanyi resumed, showing that he assimilated the punch: “Any heuristic process of a routine character—for which in the deductive sciences the Gödelian process is an example—could likewise be carried out automatically.” However, Polanyi used the same argument to dismiss the game of chess as a testbed for machine intelligence, noting: “A routine game of chess can be played automatically by a machine, and indeed, all arts can be performed automatically to the extent to which the rules of the art can be specified.”
@@
Chess, not conversation, had been Turing’s chosen field to illustrate, develop, and test machine intelligence since at least February 1946.333333“Proposed electronic calculator,” February 1946. Archives Centre, King’s College, Cambridge, AMT/C/32. On p. 16, Turing asks: “Can the machine play chess?” In his 1948 ‘Intelligent Machinery’ (op. cit., pp. 21-22), Turing had discussed a tradeoff between convenient and impressive intellectual fields for exploring machine intelligence. After discussing “various games e.g. chess,” he wrote: “Of the above possible fields the learning of languages would be the most impressive, since it is the most human of these activities.” However, he avoided language learning because it seemed “to depend rather too much on sense organs and locomotion to be feasible,” stuck with chess, and ended up describing a chess-based imitation game. Now, in October 1949, he saw chess being dismissed as unimpressive to make the case for machine intelligence because its rules could be specified.
@@
Some time later, probably around Christmas 1949, Turing read Jefferson’s Lister Oration [jefferson1949] and marked the passage quoting René Descartes (p. 1106), which starts: “Descartes made the point, and a basic one it is, that a parrot repeated only what it had been taught and only a fragment of that; it never used words to express its own thoughts.” Overall, Jefferson suggested ‘speech’ to be the distinguishing feature of human intelligence compared to other kinds of animal intelligence: “Granted that much that goes on in our heads is wordless $\ldots$ we certainly require words for conceptual thinking as well as for expression $\ldots$ It is here that there is the sudden and mysterious leap from the highest animal to man, and it is in the speech areas of the dominant hemisphere $\ldots$ that Descartes should have put the soul, the highest intellectual faculties” (p. 1109).
@@
Unlike chess, which is governed by definite rules, good performance in conversation cannot be easily specified. Turing’s 1950 choice for “the learning of languages” as the intellectual field addressed in his test can be best understood as yet another concession to Jefferson and, in this case, to Polanyi as well.
Finally, Jefferson also argued that the nervous impulse is not a purely electrical phenomenon but also a chemical one that depends on the continuity of specific physical quantities (p. 1108). It would thus be incommensurable with the activity of a digital computer. In response, Turing formulated “(7) The Argument from Continuity in the Nervous System” [turing1950, p. 451], in which he used the imitation game in its machine-machine version to neutralize that physical difference. A digital computer (a discrete-state machine) could imitate a differential analyzer (a continuous-state machine) to compute a transcendental number such as $\pi$ up to a certain digit. Turing gave this as a proof of concept that such a difference in nature disappears with the power of universal digital computing, given sufficient storage.
In summary, there is enough evidence to suggest that Turing varied the design of his imitation tests in response to the challenges posed by Hartree, Polanyi, and Jefferson. Turing’s test was a response to critics. But was it also intended as a positive proposition?
In the 1990s, Turing’s former student, close friend, and literary executor, Robin Gandy, wrote that Turing’s 1950 paper “was intended not so much as a penetrating contribution to philosophy but as propaganda.” Gandy added: “Turing thought the time had come for philosophers and mathematicians and scientists to take seriously the fact that computers were not merely calculating engines but were capable of behaviour which must be accounted as intelligent; he sought to persuade people that this was so. He wrote this paper unlike his mathematical papers quickly and with enjoyment. I can remember him reading aloud to me some of the passages always with a smile, sometimes with a giggle” [gandy1996, p. 125]. We can now explore the effect of Turing’s propaganda for machine intelligence on the other side of the North Atlantic.
TURING’S TEST AND EARLY AI
Claude Shannon visited Turing in Manchester in October 1950, [^34] Claude E. Shannon, an oral history conducted in 1982 by Robert Price. IEEE History Center, Piscataway, NJ, USA. and may have returned to the United States with an offprint of Turing’s 1950 paper, [^35] The first reprint of Turing’s ‘Computing Machinery and Intelligence’ in the US appears to be that of James R. Newman (Ed.), The World of Mathematics, vol. 4 (New York: Simon and Schuster), first published January 1, 1956. which he would cite in his “Computers and Automata” [shannon1953] published in Proc. IRE in October 1953. “Rereading Samuel Butler’s Erewhon today,” Shannon wrote, “one finds ‘The Book of the Machines’ disturbingly prophetic” (p. 1235). Butler’s novel [butler1872] envisioned a revolution of the machines against a satire of the Victorians representing the human species. It was invoked as a dystopia in June 1949, first by a Catholic priest in a letter to The Times published on the 14th (p. 5), and then in an editorial, “Mind, Machine, and Man,” of the British Medical Journal introducing Jefferson’s article on the 25th (pp. 1129-1130). Both urged scientists to disassociate themselves from Turing’s research program. Turing responded with irony, not without a point. In his writings of 1950 and 1951 he referred to Butler’s work for an image of his vision of the future of machines in nature and society [goncalves2023irony].